# Data-driven Model Predictive Control: Asymptotic Stability despite Approximation Errors exemplified in the Koopman framework

### v1

@@ 3. Asymptotic stability of MPC
A common method to ensure closed-loop stability is the use of suitably constructed terminal conditions, see, e.g., chen1998quasi ; magni2001stabilizing . However, the design of terminal ingredients can be a demanding task for nonlinear systems. A possible alternative to achieve asymptotic stability without terminal conditions relies on cost controllability, a controllability condition originally proposed in GrimMess05 ; TunaMess06 ; GrunPann10 ; Wort11 , see Wort12 for a unifying comparison, in combination with a sufficiently long prediction horizon $N$, see CoroGrun20 and KohlZeil23 for an extension to detectable stage costs. We follow the recent formulation proposed in (CoroGrun20, , Assumption 1).
Assumption 4 (Cost controllability on a set $S$).
Consider a set $S\subseteq\mathbb{R}^{n}$ containing the origin $0=0_{\mathbb{R}^{n}}$ in its interior. The system (1) with stage cost (6) is cost controllable on the set $S$, i.e., there exists a monotonically increasing and bounded sequence $(B_{N})_{N\in\mathbb{N}}=(B_{N}(S))_{N\in\mathbb{N}}$ such that, for every state $\hat{x}\in S$, there exists a control sequence $\mathbf{u}\in\mathcal{U}_{\infty}^{\varepsilon}(\hat{x})$ satisfying the growth bound [^2] The first inequality directly follows from the definition of the value function and is only stated to link (8) to cost controllability as stated in GrunPann10 .
$V_{N}(\hat{x})\leq J_{N}(\hat{x},\mathbf{u})\leq B_{N}\ell^{\star}(\hat{x})\qquad\forall\,N\in\mathbb{N},$ (8)
where $\ell^{\star}(\hat{x}):=\min_{u\in\mathbb{U}}\ell(\hat{x},u)=\|\hat{x}\|_{Q}^{2}$.
Next, we derive asymptotic stability of the origin w.r.t. the MPC closed loop relying on the surrogate (3) in the optimization step. To this end, we first show in Proposition 1 that the growth condition (8) is essentially preserved if only the approximation is sufficiently accurate, i.e., $c_{x}^{\varepsilon}$ and $c_{u}^{\varepsilon}$ in the proportional error bound of Assumption 1 are sufficiently small. Then, we conclude asymptotic stability in Theorem 1 and, finally, even prove cost controllability in Corollary 1. To this end, the combination of proportional bound and uniform Lipschitz continuity is key.
@@
Consider a set $S\subseteq\mathbb{R}^{n}$ containing the origin $0=0_{\mathbb{R}^{n}}$ in its interior and let Assumptions 1 to 3 hold. Given $\bar{N}\in\mathbb{N}$, suppose that the growth condition (8) is satisfied for all $N\in[1:\bar{N}]$ with $\mathbf{u}\in\mathcal{U}_{\bar{N}}^{\varepsilon}(\hat{x})\subset\mathcal{U}_{\bar{N}}(\hat{x})$. Then, there exists a monotonically increasing sequence $(B^{\varepsilon}_{N})_{N\in[1:\bar{N}]}$ such that
$V^{\varepsilon}_{N}(\hat{x})\leq J^{\varepsilon}_{N}(\hat{x},\mathbf{u})\leq B^{\varepsilon}_{N}\cdot\ell^{\star}(\hat{x})\qquad\forall\,N\in[1:\bar{N}].$ (9)
Moreover, $B^{\varepsilon}_{N}\rightarrow B_{N}$, $N\in[1:\bar{N}]$, for $c_{x}^{\varepsilon}\rightarrow 0$ and $c_{u}^{\varepsilon}\to 0$.
Proof. Let $\bar{\lambda}$ and $\underline{\lambda}$ denote the largest and the smallest of the (union of the) eigenvalues of the (positive-definite) weighting matrices $Q$ and $R$, respectively. For given $\bar{N}\in\mathbb{N}$, let $\hat{x}\in S$ and $\mathbf{u}\in\mathcal{U}_{\bar{N}}^{\varepsilon}(\hat{x})$ satisfy the growth bound (8). [^3] Note that $\mathbf{u}\in\mathcal{U}_{\bar{N}}^{\varepsilon}(\hat{x})$ implies, invoking Assumption 2, that $x_{\mathbf{u}}(k)\in S$ for all $k\in[0:\bar{N}]$. Hence, the growth bound (8) can be leveraged, again, for $x_{\mathbf{u}}(k)$. Then, we have [^4] We use $\pm$ to indicate that we add and subtract the following quantity.
$\underbrace{\ell(x_{\mathbf{u}}^{\varepsilon}(k),u(k))}_{=\|x_{\mathbf{u}}^{\varepsilon}(k)\mp x_{\mathbf{u}}(k)\|^{2}_{Q}+\|u(k)\|^{2}_{R}}-\ell(x_{\mathbf{u}}(k),u(k))\leq\bar{\lambda}\|x_{\mathbf{u}}^{\varepsilon}(k)-x_{\mathbf{u}}(k)\|^{2}+2\bar{\lambda}\|x_{\mathbf{u}}^{\varepsilon}(k)-x_{\mathbf{u}}(k)\|\|x_{\mathbf{u}}(k)\|.$ (10)
@@
$J^{\varepsilon}_{N}(\hat{x},\mathbf{u})$ $\leq J_{N}(\hat{x},\mathbf{u})+\bar{\lambda}\Big{(}\sum\nolimits_{k=1}^{N-1}e_{k}^{2}+2e_{k}\|x_{\mathbf{u}}(k)\|\Big{)}.$
Hence, we next derive bounds on $e_{k}^{2}$ and $e_{k}\|x_{\mathbf{u}}(k)\|$. To this end, the error $e_{k+1}:=\|x_{\mathbf{u}}^{\varepsilon}(k+1)-x_{\mathbf{u}}(k+1)\|$ at time $k+1$ is estimated using the triangle inequality by
@@
$\leq c_{x}\|x_{\mathbf{u}}^{\varepsilon}(k)\|+c_{u}\|u(k)\|+L_{f}\|x_{\mathbf{u}}^{\varepsilon}(k)-x_{\mathbf{u}}(k)\|$
$=c_{x}\|x_{\mathbf{u}}^{\varepsilon}(k)\pm x_{\mathbf{u}}(k)\|+c_{u}\|u(k)\|+L_{f}\|x_{\mathbf{u}}^{\varepsilon}(k)-x_{\mathbf{u}}(k)\|$
$\leq c_{x}\|x_{\mathbf{u}}(k)\|+c_{u}\|u(k)\|+(L_{f}+c_{x})e_{k}$
$\leq\bar{c}(\|x_{\mathbf{u}}(k)\|+\|u(k)\|)+(L_{f}+c_{x})e_{k}$ (11)
$=\bar{c}\sum\nolimits_{i=0}^{k}(L_{f}+c_{x})^{k-i}(\|x_{\mathbf{u}}(i)\|+\|u(i)\|).$ (12)
with $\bar{c}:=\max\{c_{x},c_{u}\}$. Let $d:=L_{f}+c_{x}$. Using inequality (11) and the fact that $(a+b)^{2}\leq 2a^{2}+2b^{2}$, we get
@@
$e_{k}\|x_{\mathbf{u}}(k)\|$$\leq$ $\bar{c}\sum\nolimits_{i=0}^{k-1}d^{k-1-i}\Big{(}\overbrace{\|x_{\mathbf{u}}(i)\|\|x_{\mathbf{u}}(k)\|+\|u(i)\|\|x_{\mathbf{u}}(k)\|}^{\leq\frac{1}{2}(\|x_{\mathbf{u}}(i)\|^{2}+\|u(i)\|^{2}+2\|x_{\mathbf{u}}(k)\|^{2})}\Big{)}$
@@
$\leq$ $\bar{c}^{2}\underbrace{\frac{4}{\underline{\lambda}}\sum\nolimits_{j=1}^{N-1}(2d^{2})^{j-1}B_{N-j}}_{=:c_{N}}\ell^{\star}(\hat{x}).$
@@
$\sum\nolimits_{k=1}^{N-1}e_{k}\|x_{\mathbf{u}}(k)\|$$\leq$ $\frac{\bar{c}}{2\underline{\lambda}}\sum_{k=1}^{N-1}\sum_{j=1}^{k}d^{j-1}\Big{(}\ell(x_{\mathbf{u}}(k-j),u(k-j))+2\ell^{\star}(x_{\mathbf{u}}(k))\Big{)}$
@@
$\leq$ $\bar{c}\underbrace{\frac{1}{2\underline{\lambda}}\bigg{[}\sum_{j=1}^{N-1}d^{j-1}B_{N-j}+2B_{N}\max_{j\in[1:N-1]}d^{j-1}(N-j)\bigg{]}}_{=:\bar{c}_{N}}\ell^{\star}(\hat{x}),$
with $d=L_{f}+c_{x}$ and $\bar{c}=\max\{c_{x},c_{u}\}$.
@@
$J^{\varepsilon}_{N}(\hat{x},\mathbf{u})$ $\leq\left(B_{N}+\bar{c}^{2}c_{N}+\bar{c}\bar{c}_{N}\right)\ell^{\star}(\hat{x})=:B_{N}^{\varepsilon}\ell^{\star}(\hat{x}),$
where, for each $N\in[1:\bar{N}]$, $B_{N}^{\varepsilon}\rightarrow B_{N}$ for $\bar{c}:=\max\{c_{x}^{\varepsilon},c_{u}^{\varepsilon}\}\searrow 0$ as claimed. $\square$
Note that, differently from cost controllability as stated in Assumption 4, the sequence $(B^{\varepsilon}_{N})_{N\in[1:\bar{N}]}$ derived in Proposition 1 is finite (and can only be prolonged if $\eta$ in Assumption 2 decreases such that the growth bound (8) holds despite the constraint tightening resulting from the compatibility condition $\mathbf{u}\in\mathcal{U}_{\bar{N}}^{\varepsilon}(\hat{x})$). However, that is sufficient to infer asymptotic stability of the origin w.r.t. the MPC closed loop resulting from the data-driven MPC scheme corresponding to Algorithm 1 for sufficiently small $\varepsilon$ (and, thus, proportionality constants $c_{x}^{\varepsilon}$ and $c_{u}^{\varepsilon}$ in Assumption 1). The existence of a (different) bounded sequence $(\tilde{B}^{\varepsilon}_{N})_{N\in\mathbb{N}}$ satisfying the characteristic growth bound (8) in cost controllability for the data-driven surrogate model will then be proved in Corollary 1.
@@
Consider a set $S\subseteq\mathbb{R}^{n}$ containing the origin $0=0_{\mathbb{R}^{n}}$ in its interior and let Assumptions 1 to 4 hold. Suppose that the growth condition (8) is satisfied with $\mathbf{u}\in\mathcal{U}_{N}^{\varepsilon}(\hat{x})\subset\mathcal{U}_{N}(\hat{x})$. Further, let the prediction horizon $N$ be chosen such that $\alpha=\alpha_{N}\in(0,1)$ holds with
@@
Then, there exists $\bar{c}_{x},\bar{c}_{u}>0$ such that the MPC controller of Algorithm 1 ensures asymptotic stability of the origin on the set $S$ on each (sub-)level set
$V_{c}^{-1}:=\{x\in S\mid V_{N}^{\varepsilon}(x)\leq c\}\subset\operatorname{int}(S),$ (14)
if $c_{x}^{\varepsilon}\leq\bar{c}_{x}$ and $c_{u}^{\varepsilon}\leq\bar{c}_{u}$.
Proof. First, we show a relaxed Lyapunov inequality for the surrogate system $f^{\varepsilon}$, following the proof of (bold2024data, , Theorem 10) in order to establish Inequality (15). To this end, note that the lower bound on the optimal value function can be inferred by
$V_{N}^{\varepsilon}(\hat{x})=\inf_{\mathbf{u}\in\mathcal{U}_{N}^{\varepsilon}(\hat{x})}J_{N}^{\varepsilon}(\hat{x},\mathbf{u})\geq\inf_{u\in\mathbb{U}}\ell(\hat{x},u)=\|\hat{x}\|_{Q}^{2}\geq\underline{\lambda}\|\hat{x}\|^{2},$
with $\underline{\lambda}>0$ defined as in the proof of Proposition 1. Further, the upper bound on the value function $V_{N}^{\varepsilon}(\hat{x})$ directly follows from the imposed (and essentially preserved) growth bound (9). Then, defining $\alpha^{\varepsilon}$ analogously to Formula (13) using the sequence $(B_{i}^{\varepsilon})_{i=2}^{N}$ instead and invoking $\lim_{\varepsilon\searrow 0}B_{i}^{\varepsilon}=B_{i}$ yields $\alpha^{\varepsilon}\in(\alpha,1)$ for sufficiently small $\varepsilon$. This ensures the relaxed Lyapunov inequality with Lyapunov function $V_{N}^{\varepsilon}$, $\varepsilon\in(0,\bar{\varepsilon}]$ by applying (GrunPann10, , Theorem 5.4). Hence, we established the value function $V_{N}^{\varepsilon}(\hat{x})$ as a Lyapunov function for the closed loop of the surrogate dynamics $f^{\varepsilon}$, i.e.
$V^{\varepsilon}_{N}(f^{\varepsilon}(x,\mu^{\varepsilon}_{N}(x)))\leq V^{\varepsilon}_{N}(x)-\alpha^{\varepsilon}\ell(x,\mu^{\varepsilon}_{N}(x)).$ (15)
In the following, it is shown that $V_{N}^{\varepsilon}$ is a Lyapunov function also for the system dynamics $f$. We have that
$V_{N}^{\varepsilon}(f(\hat{x},\mu_{N}^{\varepsilon}(\hat{x})))$ $=V_{N}^{\varepsilon}(f(\hat{x},\mu_{N}^{\varepsilon}(\hat{x})))\pm V_{N}^{\varepsilon}(f^{\varepsilon}(\hat{x},\mu_{N}^{\varepsilon}(\hat{x})))$
$\leq V_{N}^{\varepsilon}(\hat{x})-\alpha^{\varepsilon}\ell(\hat{x},\mu_{N}^{\varepsilon}(\hat{x}))+V_{N}^{\varepsilon}(f(\hat{x},\mu_{N}^{\varepsilon}(\hat{x})))-V_{N}^{\varepsilon}(f^{\varepsilon}(\hat{x},\mu_{N}^{\varepsilon}(\hat{x}))).$ (16)
Next, we exploit the proportional bounds of Assumption 1 to derive a bound for $V_{N}^{\varepsilon}(f(\hat{x},\mu_{N}^{\varepsilon}(\hat{x})))-V_{N}^{\varepsilon}(f^{\varepsilon}(\hat{x},\mu_{N}^{\varepsilon}(\hat{x})))$. To this end, we recall that $f^{\varepsilon}(\hat{x},\mu_{N}^{\varepsilon}(\hat{x}))=x_{\mathbf{u}^{\star}}^{\varepsilon}(1;\hat{x})$, where $\mathbf{u}^{\star}$ is the optimal solution of (OCP) with initial state $\hat{x}$, and we denote the real successor state of the system by $x^{+}:=f(\hat{x},\mu_{N}^{\varepsilon}(\hat{x}))$. We denote $\mathbf{u}^{\sharp}=(u^{\sharp}(i))_{i=0}^{N-1}$ the solution of MPC optimization starting from $\tilde{x}^{+}:=x_{\mathbf{u}^{\star}}^{\varepsilon}(1;\hat{x})$ (which is never computed in practice, but which is the sequence that is needed to define $V^{\varepsilon}_{N}(f^{\varepsilon}(\hat{x},\mu^{\varepsilon}_{N}(\hat{x})))$ in the relaxed Lyapunov inequality). Then, for the optimality of MPC, we have that
$\underbrace{V_{N}^{\varepsilon}(f(\hat{x},\mu_{N}^{\varepsilon}(\hat{x})))}_{=V_{N}^{\varepsilon}(x^{+})\leq J_{N}^{\varepsilon}(x^{+},\mathbf{u}^{\sharp})}-V_{N}^{\varepsilon}(f^{\varepsilon}(\hat{x},\mu_{N}^{\varepsilon}(\hat{x})))\leq\sum_{i=0}^{N-1}\Big{(}\ell(x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+}),u^{\sharp}(i))-\ell(x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+}),u^{\sharp}(i))\Big{)}.$ (17)
@@
$\ell(x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+}),u^{\sharp}(i))-\ell(x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+}),u^{\sharp}(i))$
$\hskip 28.45274pt=\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+})\|_{Q}^{2}-\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|_{Q}^{2}$
$\hskip 28.45274pt\leq\|Q\|\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+})-x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|\cdot\underbrace{\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+})+x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|}_{=\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+})-x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})+2x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|}$
$\hskip 28.45274pt\leq 2\|Q\|\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})-x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+})\|+\|Q\|\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})-x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+})\|^{2},$ (18)

### v2

@@ 3. Asymptotic stability of MPC
A common method to ensure closed-loop stability is the use of suitably constructed terminal conditions, see, e.g., Chen and Allgöwer (1998); Magni et al. (2001). However, the design of terminal ingredients can be a demanding task for nonlinear systems. A possible alternative to achieve asymptotic stability without terminal conditions relies on cost controllability in combination with a sufficiently long prediction horizon $N$ to ensure a relaxed Lyapunov inequality
$)V_{N}(f(\hat{x},\mu_{N}(\hat{x})))\leq V_{N}(\hat{x})-\alpha_{N}\ell(\hat{x},\mu_{N}(\hat{x}$))
with $\alpha_{N}\in(0,1]$, where $\mu_{N}$ denots the MPC control law using the original dynamics $f$ in the optimization step (2), see Worthmann (2011); Coron et al. (2020) and also Köhler et al. (2023) for an extension to detectable stage costs. Then, asymptotic stability of the MPC closed loop can be concluded using relaxed dynamic programming Grüne and Rantzer (2008). $\alpha_{N}\in(0,1]$ is called suboptimality degree (performance index) since the MPC closed-loop costs on the infinite horizon are bounded by $V_{\infty}(\hat{x})/\alpha_{N}$, e.g., a factor two for $\alpha_{N}=0.5$. The computation of the suboptimality degree $\alpha_{N}$ using cost controllability was originally proposed in Grimm et al. (2005); Tuna et al. (2006) as well as further elaborated in Grüne (2009); Grüne et al. (2010), see also Worthmann (2012) for a unifying comparison.
We follow the recent formulation of cost controllability proposed in (Coron et al., 2020, Assumption 1).
Definition 2 (Cost controllability on a set $S$).
Consider a set $S\subseteq\mathbb{R}^{n}$ containing the origin in its interior. The system (1) with stage cost (4) is cost controllable on the set $S$ if there exists a monotonically increasing and bounded sequence $(B_{N})_{N\in\mathbb{N}_{0}}=(B_{N}(S))_{N\in\mathbb{N}_{0}}$ such that for every $\hat{x}\in S$ there exists a control sequence $\mathbf{u}=\mathbf{u}(\hat{x})\in\mathcal{U}_{\infty}$ that renders the set $S$ invariant and satisfies the growth bound [^3] The first inequality follows from the definition of the value function and is only stated to link (5) to cost controllability.
[…]
for all $N\in\mathbb{N}$ with $\ell^{\star}(\hat{x}):=\min_{u\in\mathbb{U}}\ell(\hat{x},u)=\|\hat{x}\|_{Q}^{2}$.
Cost controllability links controllability to the performance measure $\ell$ and, thus, ensures that the the system can be controlled to the origin sufficiently fast if only the prediction horizon $N$ is chosen large enough.
Next, we derive asymptotic stability of the origin w.r.t. the MPC closed loop relying on the surrogate (3) in the optimization step. To this end, we first show in Proposition 1 that the growth condition (5) is essentially preserved if only the approximation is sufficiently accurate thanks to (P-bound) and (U-bound) and the assumed Lipschitz continuity of the surrogate model. Then, we conclude asymptotic stability in Theorem 1 and, finally, even show cost controllability in Corollary 1.
@@
Suppose that system (1) is cost controllable on a set $S$. Further, let $\bar{N}\in\mathbb{N}$ be given and the Assumptions 1 and 2 hold on a set $\Omega$ satisfying $S\oplus\mathcal{B}_{r(\bar{N})\eta^{\varepsilon}}(0)\subseteq\Omega$ with $r(\bar{N}):=\sum_{i=0}^{\bar{N}-1}\bar{L}^{i}$ for all $\varepsilon\in(0,\bar{\varepsilon}]$. Then, for the surrogate model (3), the growth bound (5) is satisfied on $S$ for all $N\in[1:\bar{N}]$ uniformly in $\varepsilon$, i.e., there exists a monotonically increasing sequence $(B_{N}^{\varepsilon})_{N\in[1:\bar{N}]}$ parametrized in $\varepsilon$ such that, for each pair $(\hat{x},N)\in S\times[1:\bar{N}]$, there exists $\mathbf{u}=\mathbf{u}(\hat{x})\in\mathcal{U}_{N}$ satisfying
[…]
Moreover, $\lim_{\varepsilon\searrow 0}B^{\varepsilon}_{N}=B_{N}$ for all $N\in[1:\bar{N}]$.
The statement holds also upon switching the roles of $f$ and $f^{\varepsilon}$, i.e. if the cost controllability condition (5) holds for the surrogate model $f^{\varepsilon}$, then the growth bound (6) holds for the original system dynamics $f$ for all $N\in[1:\bar{N}]$.
Proposition 1 shows that the growth bound (5) characterizing cost controllability of the system (1) is preserved for the surrogate model (3) up to a maximal horizon length $\bar{N}$. The restriction to an arbitrary, but fixed maximal horizon is necessary to apply the proposed proof technique in view of the condition $S\oplus\mathcal{B}_{r(\bar{N})\eta^{\varepsilon}}(0)\subseteq\Omega$ for all $\varepsilon\in(0,\bar{\varepsilon}]$. Otherwise, one has either to enlarge the set $\Omega$, on which Assumptions 1 and 2 hold or reduce $\bar{\varepsilon}$.
Proof. For given $N\in\mathbb{N}$, let $\hat{x}\in S$ and $\mathbf{u}\in\mathcal{U}_{N}$ satisfy the growth bound (5).
First, we show $x_{\mathbf{u}}^{\varepsilon}(k)\in\Omega$, $k\in[1:N]$. Note that cost controllability yields $x_{\mathbf{u}}(k)\in S$, $k\in[1:\bar{N}]$. We show $\|x_{\mathbf{u}}^{\varepsilon}(k)-x_{\mathbf{u}}(k)\|\leq r(k)\eta^{\varepsilon}=\sum_{i=0}^{k-1}L_{f^{\varepsilon}}^{i}\eta^{\varepsilon}$ by induction. For $k=1$, we exploit (U-bound) to derive
$\|x_{\mathbf{u}}^{\varepsilon}(1)-x_{\mathbf{u}}(1)\|=\|f(\hat{x},u(0))-f^{\varepsilon}(\hat{x},u(0))\|\leq\eta^{\varepsilon}$
Now, let the assertion hold for $k$. Then, we have
$\|x_{\mathbf{u}}^{\varepsilon}(k+1)-x_{\mathbf{u}}(k+1)\|$
$=$ $\|f^{\varepsilon}(x_{\mathbf{u}}^{\varepsilon}(k),u(k))\mp f^{\varepsilon}(x_{\mathbf{u}}(k),u(k))-f(x_{\mathbf{u}}(k),u(k))\|$
$\leq$ $\eta^{\varepsilon}+L_{f^{\varepsilon}}\|x_{\mathbf{u}}^{\varepsilon}(k)-x_{\mathbf{u}}(k)\|\leq\sum\nolimits_{i=0}^{k}L_{f^{\varepsilon}}^{i}\eta^{\varepsilon}.$
Hence, we have $x_{\mathbf{u}}^{\varepsilon}(k)\in S\oplus\mathcal{B}_{r(k)\eta^{\varepsilon}}(0)\subseteq S\oplus\mathcal{B}_{r(\bar{N})\eta^{\varepsilon}}(0)\subseteq\Omega$ for all $k\in[0:\bar{N}]$, so that Assumptions 1 and 2 hold.
Next, study the difference between $J_{N}(\hat{x},\mathbf{u})$ and $J_{N}^{\varepsilon}(\hat{x},\mathbf{u})$ using the notation $\underline{\lambda}:=\operatorname{min}\{\lambda_{\min}(Q),\lambda_{\min}(R)\}$ and $\bar{\lambda}:=\lambda_{\max}(Q)$. We have
$\underbrace{\ell(x_{\mathbf{u}}^{\varepsilon}(k),u(k))}_{=\|x_{\mathbf{u}}^{\varepsilon}(k)\mp x_{\mathbf{u}}(k)\|^{2}_{Q}+\|u(k)\|^{2}_{R}}-\ell(x_{\mathbf{u}}(k),u(k))$
$\leq$ $\bar{\lambda}\|x_{\mathbf{u}}^{\varepsilon}(k)-x_{\mathbf{u}}(k)\|^{2}+2\bar{\lambda}\|x_{\mathbf{u}}^{\varepsilon}(k)-x_{\mathbf{u}}(k)\|\|x_{\mathbf{u}}(k)\|.$
@@
$J^{\varepsilon}_{N}(\hat{x},\mathbf{u})$ $\leq J_{N}(\hat{x},\mathbf{u})+\bar{\lambda}\Big[\sum\nolimits_{k=1}^{N-1}e_{k}^{2}+2e_{k}\|x_{\mathbf{u}}(k)\|\Big].$
Next, we derive bounds on $e_{k}^{2}$ and $e_{k}\|x_{\mathbf{u}}(k)\|$. To this end, the error $e_{k+1}:=\|x_{\mathbf{u}}^{\varepsilon}(k+1)-x_{\mathbf{u}}(k+1)\|$ at time $k+1$ is estimated using the triangle inequality by
@@
[…]
with $\bar{c}:=\max\{c_{x}^{\varepsilon},c_{u}^{\varepsilon}\}$. Let $d:=L_{f}+c_{x}^{\varepsilon}$. Using inequality (8) and the fact that $(a+b)^{2}\leq 2a^{2}+2b^{2}$, we get
@@
$e_{k}\|x_{\mathbf{u}}(k)\|$
$\leq$ $\bar{c}\sum\nolimits_{i=0}^{k-1}d^{k-1-i}\Big(\overbrace{\|x_{\mathbf{u}}(i)\|\|x_{\mathbf{u}}(k)\|+\|u(i)\|\|x_{\mathbf{u}}(k)\|}^{\leq\frac{1}{2}(\|x_{\mathbf{u}}(i)\|^{2}+\|u(i)\|^{2}+2\|x_{\mathbf{u}}(k)\|^{2})}\Big)$
@@
[…]
@@
$\sum\nolimits_{k=1}^{N-1}e_{k}\|x_{\mathbf{u}}(k)\|$
$\leq$ $\frac{\bar{c}}{2\underline{\lambda}}\sum_{k=1}^{N-1}\sum_{j=1}^{k}d^{j-1}\Big(\ell(x_{\mathbf{u}}(k-j),u(k-j))+2\ell^{\star}(x_{\mathbf{u}}(k))\Big)$
@@
[…]
with $d=L_{f}+c_{x}^{\varepsilon}$ and $\bar{c}=\max\{c_{x}^{\varepsilon},c_{u}^{\varepsilon}\}$.
@@
[…]
where, for each $N\in[1:\bar{N}]$, $B_{N}^{\varepsilon}\rightarrow B_{N}$ for $\bar{c}:=\max\{c_{x}^{\varepsilon},c_{u}^{\varepsilon}\}\searrow 0$ as claimed.
Finally, note that the proof is only based on the bounds on the difference between $x_{\mathbf{u}}(k)$ and $x_{\mathbf{u}}^{\varepsilon}(k)$. Hence, the same reasoning can be applied to derive a bound on $J_{N}(\hat{x},\mathbf{u})$ starting from $J_{N}^{\varepsilon}(\hat{x},\mathbf{u})$. $\square$
Note that, differently from cost controllability as stated in Definition 2, the sequence $(B^{\varepsilon}_{N})_{N\in[1:\bar{N}]}$ derived in Proposition 1 is finite. In particular, if $\max\{c_{x}^{\varepsilon},c_{u}^{\varepsilon}\}>0$, the terms $B_{N}^{\varepsilon}$ defined in (10) grow unboundedly for $N\to\infty$. However, a finite sequence is sufficient to infer asymptotic stability of the origin w.r.t. the MPC closed loop resulting from Algorithm 1 for sufficiently small $\varepsilon$ (and, thus, sufficiently small $c_{x}^{\varepsilon}$ and $c_{u}^{\varepsilon}$ in Assumption 1). The existence of a bounded sequence $(B^{\varepsilon}_{N})_{N\in\mathbb{N}}$ of infinite length satisfying the characteristic growth bound (5) of cost controllability for the data-driven surrogate model will be proved in Corollary 1 leveraging the then established asymptotic stability.
The results of Proposition 1 are similar to (Bold et al., 2025a, Proposition 9), but a tighter bound on the difference between $B_{N}$ and $B_{N}^{\varepsilon}$ is provided and it is clarified that for a given $\varepsilon>0$, the sequence $(B_{N}^{\varepsilon})_{N\in[1:\bar{N}]}$ can only be derived up to some horizon $\bar{N}$. In addition, the relation between the sets $S$ and $\Omega$ has been clarified. Moreover, it is shown that the statement is symmetric and holds upon switching the role of the system and the surrogate model. This property, which was not studied in Bold et al. (2025a), can be particularly useful in practical applications, in which the real dynamics is unknown and cost controllability can only be checked on the surrogate model.
@@
Let the assumptions of Proposition 1 hold with an optimization horizon $N=\bar{N}$ such that $\alpha_{N}\in(0,1]$ holds, where $\alpha_{N}$ is given by [^4] The condition $\alpha_{N}\in(0,1]$ is always satisfied for sufficiently large $N$ according to Grüne et al. (2010) if the system (1) is cost controllable with a bounded sequence $(B_{N})_{N\in\mathbb{N}_{0}}$.
@@
Then, there exists $\varepsilon_{0}\in(0,\bar{\varepsilon}]$ such that the MPC controller of Algorithm 1 ensures asymptotic stability of the origin on each sublevel set $V_{N}^{-\varepsilon}(c):=\{x\in\mathbb{R}^{n}\mid V_{N}^{\varepsilon}(x)\leq c\}$ satisfying the set inclusion $V_{N}^{-\varepsilon}(c)\subseteq S$ for $\varepsilon\in(0,\varepsilon_{0})$.
Proof. First, we show the relaxed Lyapunov inequality
$V^{\varepsilon}_{N}(f^{\varepsilon}(x,\mu^{\varepsilon}_{N}(x)))\leq V^{\varepsilon}_{N}(x)-\alpha^{\varepsilon}_{N}\ell(x,\mu^{\varepsilon}_{N}(x))$ (12)
for the surrogate model $f^{\varepsilon}$ following the proof of (Bold et al., 2025a, Theorem 10). First, we infer a lower bound on the optimal value function by
$V_{N}^{\varepsilon}(\hat{x})=\inf_{\mathbf{u}\in\mathcal{U}_{N}}J_{N}^{\varepsilon}(\hat{x},\mathbf{u})\geq\inf_{u\in\mathbb{U}}\ell(\hat{x},u)=\|\hat{x}\|_{Q}^{2}\geq\underline{\lambda}\|\hat{x}\|^{2},$
[…]
In the following, we derive novel bounds to show that $V_{N}^{\varepsilon}$ is a Lyapunov function also for the system dynamics $f$. We have that
[…]
$\leq$ $\sum_{i=0}^{N-1}\Big(\ell(x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+}),u^{\sharp}(i))-\ell(x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+}),u^{\sharp}(i))\Big).$
@@
$\ell(x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+}),u^{\sharp}(i))-\ell(x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+}),u^{\sharp}(i))$ (15)
$=$ $\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+})\|_{Q}^{2}-\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|_{Q}^{2}$
$\leq$ $\|Q\|\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+})-x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|\cdot\underbrace{\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+})+x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|}_{=\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+})-x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})+2x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|}$
$\leq$ $2\|Q\|\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})-x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+})\|$
$+\|Q\|\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})-x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+})\|^{2},$